# Optimizing Visible Light Communication Efficiency Through Reinforcement Learning-Based NOMA-CSK Integration

### v1

@@ Abstract
In this paper, we explore the use of Non-Orthogonal Multiple Access (NOMA) and Color Shift Keying (CSK) for Visible Light Communication (VLC) systems. VLC is a wireless communication technology that uses visible light as the carrier signal to transmit information. It has several advantages over traditional radio frequency communication, including higher bandwidth, lower interference, and greater security. We first provide an introduction to NOMA and CSK and explain how they can be applied to VLC systems. NOMA is a technique that allows multiple users to share the same frequency channel by allocating different power levels to each user. This enables more users to connect to a single VLC transmitter simultaneously, thereby improving system capacity and spectral efficiency. CSK, on the other hand, is a modulation technique that uses different colors of light to represent digital information. By changing the color of the transmitted signal, information can be encoded and decoded at the receiver. Next, we discuss how NOMA and CSK can be combined in VLC systems by using different power levels to represent different users. This allows for more efficient use of the frequency spectrum, as multiple users can share the same channel at the same time. Additionally, we examine the potential benefits of using NOMA and CSK together in VLC systems to increase data rate. Finally, we discuss how reinforcement learning, a machine learning technique used to train agents to make decisions based on environmental feedback, can be used to optimize NOMA-CSK-VLC networks by allowing agents to learn and adapt to changing network conditions. Overall, our paper provides insights into the benefits of combining NOMA and CSK for VLC systems, highlighting the potential for improving communication efficiency and performance.
@@ 1 Introduction
With the development and widespread use of wireless communication systems, the number of wireless devices and data transfer requirements are increasing. However, the current frequency band of 6 GHz and below has become too crowded to provide efficient communication [1]. Therefore, it is suggested to use wide unlicensed spectrum of the visible light in wireless communication systems [2]. Visible light communication (VLC) systems offer a number of benefits, including energy savings due to the ability to carry out lighting and communication simultaneously, security, privacy, and resistance to electromagnetic radiation [3, 4].
@@
VLC systems usually utilize light emitting diodes (LEDs) as transmitters, while photodetectors (PDs) are favored as receivers due to their quick response times. The modulation speeds of LEDs are a crucial element in communication performance. While phosphor-coated white LEDs have modulation speeds of approximately 10 MHz, multicolored LEDs, such as Red, Green Blue, (RGB) and Quadrature LED (QLED), can be modulated at speeds of almost 100 MHz. Therefore, compared to white LEDs with phosphor coating, communication via multicolored LEDs may theoretically be possible at higher speeds.
@@
In wireless communication systems, modulation type, which is one of the factors affecting communication performance, is as important as date rates. Various modulation methods such as On-Off Keying (OOK) [5], Pulse Width Modulation (PWM) [6], Color Shift Keying (CSK) [2] can be used as modulation type in VLC systems. The newest of these, CSK, is a modulation method published in the IEEE 802.15.7 standard and used in VLC systems with multi-colored LED transmitters [2]. The CSK modulation method is a modulation method in which the color of the light emitted from the transmitter is changed according to the information signal in multicolor LED VLC systems. On the receiver side, there are color filters. Signals obtained from color filters are used in demodulation. For RGB LED emitter systems, the IEEE 802.15.7 standard defines 4-CSK, 8-CSK, and 16-CSK.
@@
CSK has several advantages in VLC systems. Firstly, it maintains a constant power envelope of the transmitted signal, which reduces the possibility of health issues caused by fluctuations in light intensity. Secondly, it provides color lights with a stable temperature. Thirdly, with rapid switching, the flicker effect is minimized. Fourthly, even at slow switching speeds, high-level modulation is possible, which increases the capacity [2]. Additionally, CSK does not require additional DC power for data transmission, giving it an edge over other modulation schemes utilized in VLC. This allows for an increase in the signal-to-noise ratio, and efficient use of the entire transmission spectrum is possible.
CSK modulation involves instantaneously changing the color of light to transmit data. As a result, lighting characteristics such as Correlated Color Temperature (CCT) [7], Color Rendering Index (CRI) [8], and illuminance [8] can also change during communication. CCT refers to the temperature of the light in Kelvin in the given environment, while CRI is a measure of how well the light source can accurately reflect the colors of objects in the environment. CRI is typically represented as a value between 0 and 100, with higher values indicating more accurate color representation.
@@
Besides CSK modulation, multiple access methods in VLC have also been extensively studied in the literature. Both orthogonal and non-orthogonal multiple access methods (OMA and NOMA) are available in VLC. Power domain NOMA (PD-NOMA), which is a non-orthogonal multiple access method, has been proposed as a promising and bandwidth-efficient solution for multiple access in indoor VLC networks [9, 10]. NOMA works by superimposing user signals in the power domain, where a separate power level is assigned to each signal based on the user’s channel conditions. In NOMA, all users have continuous access to full bandwidth. The fundamental tenet of NOMA is to give users with weak channels more power than those with strong channels. To achieve this, the user with the highest power can detect the information signal despite the interference caused by other users, while other users perform successive interference cancellation (SIC) to decode their own signals [11]. It has been demonstrated in [12] that NOMA can achieve higher data rates than Orthogonal Frequency Division Multiple Access (OFDMA) in VLC systems .Similarly, in [13] researchers examined the performance of indoor VLC systems using NOMA and discovered a substantial enhancement in system capacity due to the implementation of NOMA. In [14], NOMA techniques were compared with classical OFDMA techniques, and it was found that some of the advantages of NOMA include better quality of service, less interferences, and the ability to provide multiple access to a larger number of users.
@@
Power allocation is a crucial aspect of NOMA VLC systems in determining the achievable capacity [15]. The gain ratio power allocation (GRPA) method was proposed in [9] for a scenario with two transmitters and three receivers, and it showed better sum-rate and Bit Error Rate (BER) than static power allocation in the best-case scenario. In [16], the normalized gain difference power allocation (NGDPA) method was introduced and compared to GRPA, resulting in 29.1% higher total capacity in a 3-user 2x2 MIMO VLC system. An optimization algorithm was proposed in [17] to maximize overall data rate and compared to GRPA. However, fairness and Quality of Service (QoS) were not considered. In [18], a convex decoder was used to optimize the total data rate and BER under a certain light intensity constraint, but users’ data rate and fairness were not considered. The VOOK method with GRPA was used in [5] for power allocation in NOMA-VLC systems with variable illumination levels. However, VOOK is inefficient as the total capacity decreases in proportion to brightness. In [19], a particle swarm optimization (PSO)-based power allocation optimization algorithm for NOMA VLC multicellular networks was proposed, but the level of fairness achieved was less than 0.6, which is suboptimal for QoS. When distributing LED transmitter power to users in VLC systems, factors such as fairness, illumination level, and user data rate should be taken into account to provide eye comfort and better QoS.
@@
Recent researches have explored the use of reinforcement learning (RL) methods to optimize communication systems. Using deep reinforcement learning (DRL) as the method, Xiao et al. [20] investigated power allocation for non-orthogonal multiple access (NOMA) against smart jammers in radio frequency (RF) systems, resulting in an increased sum-rate. He et al. [21], on the other hand, studied power distribution in multi-carrier RF systems with NOMA and observed an overall system performance improvement. Zhang et al. [22] employed DRL to investigate energy-efficient resource allocation, while Zhang et al. [23] demonstrated a 32.9% higher efficiency compared to the ALOHA method. Wang et al. [24] proposed the DRL-JRM (Deep Reinforcement Learning-Joint Resource Management) scheme for RF multi-carrier systems in NOMA, which showed superior system efficiency and resistance to interference, especially in the presence of large numbers of users and strong intercellular interference. Bariah et al. [25] investigated Deep Q-Learning-Based resource allocation to increase the energy efficiency and sum-rate for NOMA VLC systems, showing its superiority over the results obtained by genetic algorithm and differential evolution. Finally, Al et al. [26] investigated Q-Learning-Based resource allocation and LED emitter angle optimization to increase sum-rate for NOMA VLC systems.
In this study, a PD-NOMA scheme was implemented in a VLC system using CSK modulation. Reinforcement learning (RL) with a Soft Actor-Critic (SAC) agent [27] was trained to maximize fairness in power allocation. The agent was trained to be applicable not only to a specific indoor environment, but to other indoor environments as well. The results showed that the proposed RL-based PD-NOMA CSK VLC system outperformed other methods in terms of fairness, sum-rate and bit error rate. Moreover, the trained agent was able to achieve the same performance in different indoor environments for the CSK VLC system without the need for retraining.
@@ 3 System Model
For the simulation purposes, an office space measuring 5 meters in length, 5 meters in width, and 3 meters in height will be utilized. The OSRAM LZ4-00MA00 multicolor LED, which is commercially available, has been selected for this particular application. The assumption made is that the emitted lighting powers for the red (R), green (G), and blue (B) colors are equivalent.
LEDs can be represented in various ways for simulation purposes, and the H model proposed in [30] closely approximates the results obtained from experimental studies. The H model parameters for a specific LED can be found in Table 1 [31], and these parameters are utilized to compute the values for Eq.2, Eq.3 and Eq.4.
$g\left(\lambda,\lambda_{p},\Delta\lambda\right)=\exp{\left[-\frac{\left(\lambda-\lambda_{p}\right)^{2}}{\mathrm{\Delta}\lambda^{2}}\right]}$ (2)
$\Delta\lambda=\left\{\begin{matrix}\Delta\lambda_{1},&\lambda<\lambda_{p}\\ \Delta\lambda_{2},&\lambda\geq\lambda_{p}\\ \end{matrix}\right.$ (3)
$PSD_{i}\left(\lambda\right)=\frac{g\left(\lambda,\lambda_{p},\Delta\lambda\right)+k_{1}g^{k_{2}}\left(\lambda,\lambda_{p},\Delta\lambda\right)}{1+k_{1}}$ (4)
Here, $\lambda_{p}$ is the peak wavelength in the $PSD_{i}$ while $i$ indicating the color, $\Delta\lambda_{1}$ and $\Delta\lambda_{2}$ are the left and right half spectral widths, respectively, and $k_{1}$ and $k_{2}$ are characteristic parameters of the PSD shape. The $PSD\left(\lambda\right)$ for each color obtained from this H model is shown in Fig.0(a).Thirty-six of these LEDs were placed on the ceiling of the office, at a distance of 95 cm from the side walls in 4 square luminaires with a distance of 3 cm between each LED, as in Fig.0(b). X and Y denote the coordinates of the bottom-left position referenced room.
@@
[…]
In the context provided, $i$ and $j$ correspond to the colors of the LED at the transmitter and the photodetector at the receiver, respectively. The values of $i$ and $j$ range from $1$ to $3$, representing the colors red, green, and blue. The variables $\lambda_{l}^{i}$ and $\lambda_{h}^{i}$ denote the lower and upper wavelengths of the spectrum for the LED color $i$, respectively. The variable $F_{j}$ represents the gain associated with the $j$-colored optical filter. The variable $S_{i}$ denotes the CSK symbol value for color $i$. The matrix $\boldsymbol{p}$ is a $3\times 3$ channel matrix, where the $(i,j)$-th element is determined by Eq. 6.
$\boldsymbol{p}=\left[\begin{matrix}p_{1,1}&p_{2,1}&p_{3,1}\\ p_{1,2}&p_{2,2}&p_{3,2}\\ p_{1,3}&p_{2,3}&p_{3,3}\\ \end{matrix}\right]$ (6)
@@
$\operatorname{SINR}=\frac{\sum\limits_{i}\sum\limits_{j}p_{i,j}\delta_{ij}}{\sum\boldsymbol{p}-\sum\limits_{i}\sum\limits_{j}p_{i,j}\delta_{ij}+N_{0}}$ (7)
Here, $\delta_{ij}$ is Kronecker delta function and $N_{0}$ is the power of AWGN noise respectively. Capacity calculation over SINR can be made done according to Eq.8. After the CSK demodulation is performed on the received signal, the throughput can be calculated using Eq.9.
$C=Blog_{2}\left({1+SINR}\right)$ (8)
$T=\ C\left(1-BER\right)$ (9)
The open-source program shared in [32] can be used to calculate the CRI, CCT, and illumination parameters mentioned in the introduction. The irradiance data corresponding to the LED’s wavelength is input into the program, which then uses it to calculate the illumination parameters. The CRI, CCT, and illumination values should be calculated over the long-term average of irradiance value because the brightness of the light emitted from the R, G, and B colors varies with communication.
4 Simulations
The variables associated with the VLC channel and their corresponding values, along with the simulation parameters, are provided in Table 2. We built a NOMA-CSK communication system with two receivers ($U_{1}$ and $U_{2}$). Both receivers’ PDs are positioned with their normal surfaces perpendicular to the ceiling.

### v2

# NOMA-CSK Integrated VLC System with Reinforcement Learning-Based Multi-Objective Power Allocation

@@ Abstract
This paper introduces a novel framework that synergistically combines Non-Orthogonal Multiple Access (NOMA) with Color Shift Keying (CSK) modulation to substantially boost spectral efficiency in Visible Light Communication (VLC) systems. A key challenge in the proposed NOMA-CSK architecture is managing the complex power allocation process, especially under cross-color interference caused by spectral overlap among LEDs and the limitations of optical filters. To overcome this, we develop an intelligent power allocation strategy powered by a Soft Actor-Critic (SAC) reinforcement learning agent. Trained in a simulated indoor environment, the SAC agent dynamically distributes power among users with diverse channel conditions while balancing multiple performance objectives. Simulation results show that our SAC-based method significantly outperforms traditional approaches such as Gain Ratio Power Allocation (GRPA) and Normalized Gain Difference Power Allocation (NGDPA), achieving superior fairness, higher overall throughput, and reduced bit error rates—even under a challenging 10 dB SNR. Notably, the trained agent demonstrates strong generalization capabilities, maintaining optimal performance in unseen environments without requiring retraining. Overall, this work makes two major contributions: it presents a pioneering NOMA-CSK VLC system design and delivers a robust, adaptive power allocation solution critical for real-world applications.
@@ 1 Introduction
With the development and widespread use of wireless communication systems, the number of wireless devices and data transfer requirements are increasing. However, the current frequency band of 6 GHz and below has become too crowded to provide efficient communication [1]. Therefore, it is suggested to use the wide unlicensed spectrum of visible light in wireless communication systems [2]. Visible Light Communication (VLC) systems offer a number of benefits, including energy savings due to the ability to carry out lighting and communication simultaneously, security, privacy, and resistance to electromagnetic radiation [3, 4].
@@
VLC systems usually utilize light-emitting diodes (LEDs) as transmitters, while photodetectors (PDs) are favored as receivers due to their quick response times. The modulation speeds of LEDs are a crucial element in communication performance. While phosphor-coated white LEDs have modulation speeds of approximately 10 MHz, multicolored LEDs, such as Red, Green, Blue (RGB) and Quadrature LED (QLED), can be modulated at speeds up to 100 MHz. Therefore, compared to white LEDs with phosphor coating, communication via multicolored LEDs may theoretically be possible at higher speeds.
@@
In wireless communication systems, the modulation type, which is one of the factors affecting communication performance, is as important as data rates. Various modulation methods such as On-Off Keying (OOK) [5], Pulse Width Modulation (PWM) [6], and Color Shift Keying (CSK) [2] can be used as modulation types in VLC systems. The newest of these, CSK, is a modulation method published in the IEEE 802.15.7 standard and used in VLC systems with multi-colored LED transmitters [2]. The CSK modulation method is a modulation method in which the color of the light emitted from the transmitter is changed according to the information signal in multicolor LED VLC systems. On the receiver side, there are color filters. Signals obtained from color filters are used in demodulation. For RGB LED emitter systems, the IEEE 802.15.7 standard defines 4-CSK, 8-CSK, and 16-CSK.
@@
CSK has several advantages in VLC systems. Firstly, it maintains a constant power envelope of the transmitted signal, which reduces the possibility of health issues caused by fluctuations in light intensity. Secondly, it provides color lights with a stable temperature. Thirdly, with rapid switching, the flicker effect is minimized. Fourthly, even at slow switching speeds, high-level modulation is possible, which increases the capacity [2]. Additionally, CSK does not require additional DC power for data transmission, giving it an edge over other modulation schemes utilized in VLC. This allows for an increase in the signal-to-noise ratio, and efficient use of the entire transmission spectrum is possible. CSK modulation involves instantaneously changing the color of light to transmit data. As a result, illuminance [7, 8] can change during communication.
@@
Besides CSK modulation, multiple access methods in VLC have also been extensively studied in the literature. Both orthogonal and non-orthogonal multiple access methods (OMA and NOMA) are available in VLC. Power domain NOMA (PD-NOMA), which is a non-orthogonal multiple access method, has been proposed as a promising and bandwidth-efficient solution for multiple access in indoor VLC networks [9, 10]. NOMA works by superimposing user signals in the power domain, where a separate power level is assigned to each signal based on the user’s channel conditions. In NOMA, all users have continuous access to the full bandwidth. The fundamental tenet of NOMA is to give users with weak channels more power than those with strong channels. To achieve this, the user with the highest power can detect the information signal despite the interference caused by other users, while other users perform successive interference cancellation (SIC) to decode their own signals [11]. It has been demonstrated in [12] that NOMA can achieve higher data rates than Orthogonal Frequency Division Multiple Access (OFDMA) in VLC systems. Similarly, in [13], researchers examined the performance of indoor VLC systems using NOMA and discovered a substantial enhancement in system capacity due to the implementation of NOMA. In [14], NOMA techniques were compared with classical OFDMA techniques, and it was found that some of the advantages of NOMA include better quality of service, less interference, and the ability to provide multiple access to a larger number of users.
@@
Power allocation is a critical challenge in NOMA-VLC systems as it directly determines the achievable capacity and user fairness [15, 16]. The gain ratio power allocation (GRPA) method was proposed in [9] for a scenario with two transmitters and three receivers, and it showed better sum-rate and Bit Error Rate (BER) than static power allocation in the best-case scenario. In [17], the normalized gain difference power allocation (NGDPA) method was introduced and compared to GRPA, resulting in a 29.1% higher total capacity in a 3-user 2x2 MIMO VLC system. An optimization algorithm was proposed in [18] to maximize the overall data rate and compared to GRPA. However, fairness and Quality of Service (QoS) were not considered. In [19], a convex decoder was used to optimize the total data rate and BER under a certain light intensity constraint, but users’ data rate and fairness were not considered. The VOOK method with GRPA was used in [5] for power allocation in NOMA-VLC systems with variable illumination levels. However, VOOK is inefficient as the total capacity decreases in proportion to brightness. In [20], a particle swarm optimization (PSO)-based power allocation optimization algorithm for NOMA VLC multicellular networks was proposed, but the level of fairness achieved was less than 0.6, which is suboptimal for QoS. When distributing LED transmitter power to users in VLC systems, factors such as fairness, illumination level, and user data rate should be taken into account to provide eye comfort and better QoS.
@@
Recent research has explored the use of reinforcement learning (RL) methods to optimize communication systems. Using deep reinforcement learning (DRL) as the method, Xiao et al. [21] investigated power allocation for non-orthogonal multiple access (NOMA) against smart jammers in radio frequency (RF) systems, resulting in an increased sum-rate. He et al. [22], on the other hand, studied power distribution in multi-carrier RF systems with NOMA and observed an overall system performance improvement. Zhang et al. [23] employed DRL to investigate energy-efficient resource allocation, while Zhang et al. [24] demonstrated a 32.9% higher efficiency compared to the ALOHA method. Wang et al. [25] proposed the DRL-JRM (Deep Reinforcement Learning-Joint Resource Management) scheme for RF multi-carrier systems in NOMA, which showed superior system efficiency and resistance to interference, especially in the presence of large numbers of users and strong intercellular interference. Bariah et al. [26] investigated Deep Q-Learning-Based resource allocation to increase the energy efficiency and sum-rate for NOMA VLC systems, showing its superiority over the results obtained by genetic algorithm and differential evolution. Al Hammadi et al. [27] proposed a Deep Q-Learning-based approach for resource allocation in NOMA Visible Light Communications (VLC), aiming to enhance the system’s efficiency and sum-rate. Finally, Al et al. [27] investigated Q-Learning-Based resource allocation and LED emitter angle optimization to increase sum-rate for NOMA VLC systems.
While NOMA and CSK have been investigated independently for VLC, their integration to leverage the benefits of both power-domain multiplexing and color-based modulation remains an unexplored area. This combination presents a unique opportunity to maximize user capacity and throughput, but it also introduces a more complex resource allocation problem. Existing power allocation schemes like GRPA and NGDPA were not designed for the specific interference characteristics of a NOMA-CSK system. It is important to note that these conventional methods were primarily designed for NOMA systems where interference could be modeled more simply, and not for the complex cross-color interference inherent in our proposed NOMA-CSK architecture. Particularly, the complex interference patterns arising from the spectral overlap of LEDs and non-ideal optical filters lead to significant cross-color interference. This complex environment makes it challenging for traditional analytical models to find an optimal and fair power allocation, necessitating an adaptive, data-driven approach. While traditional optimization methods can find solutions for well-defined, static channel models, they struggle to adapt to the dynamic and stochastic nature of real-world VLC environments with complex, non-linear cross-color interference. Reinforcement learning, particularly the SAC agent, offers a powerful alternative by learning a robust control policy directly from environmental interactions, enabling it to generalize to unseen channel conditions without requiring an explicit analytical model of the complex interference.
To address this research gap, this paper makes a twofold contribution. First, we propose and model a novel PD-NOMA VLC system that utilizes CSK modulation (NOMA-CSK). Second, we solve the inherent power allocation challenge within this new system by developing a fairness-driven SAC reinforcement learning agent. Our key contributions are:
  • A Novel NOMA-CSK System Architecture: We design and model a VLC system that combines power-domain NOMA with CSK, establishing a framework for efficient multi-user communication.
  • A Fairness-Centric Multi Ojective RL Power Allocation Solution: Within this new system, we design and train a SAC agent with a reward mechanism that focuses on Jain’s fairness index,ensuring equitable resource distribution, along with the user rates, bit error rates, power efficiency and stability.
  • Superior Performance and Robustness: We demonstrate through extensive simulations that our RL-based approach significantly outperforms traditional benchmarks in sum-rate, BER, and fairness. Crucially, we prove the agent’s ability to generalize, as it maintains high performance in unseen environments without retraining, underscoring its practical value.
This work lays the foundation for NOMA-CSK as a viable technology and provides the critical intelligent mechanism needed for its efficient operation.
@@ 2 VLC Channel
It is important to note that this study primarily considers the Line-of-Sight (LoS) transmission model, which constitutes the dominant signal component in most indoor VLC scenarios. While this is a common and valid simplification for tractable analysis, real-world environments also include Non-Line-of-Sight (NLoS) components arising from reflections. These reflections could alter the channel characteristics and interference profile. The investigation of NLoS effects on the performance of the NOMA-CSK system and the adaptability of the SAC agent to such conditions remains a valuable direction for future work.
@@ 3 System Model
The simulations are conducted in a virtual office space measuring 5 meters in length, 5 meters in width, and 3 meters in height. The OSRAM LZ4-00MA00 multicolor LED, which is commercially available, has been selected for this particular application. We assumed that the emitted lighting powers for the red (R), green (G), and blue (B) colors are equivalent.
LEDs can be represented in various ways for simulation purposes, and the H model proposed in [30] closely approximates the results obtained from experimental studies. The H model parameters for LZ4-00MA00 LED can be found in Table 1 [31], and these parameters are utilized to compute the Power Spectral Density (PSD) (Eq.2) of the illuminating LED by using,Eq.3 and Eq.4.
$g(\lambda,\lambda_{p},\Delta\lambda)=\exp{\left[-\frac{(\lambda-\lambda_{p})^{2}}{\Delta\lambda^{2}}\right]}$ (2)
$\Delta\lambda=\begin{cases}\Delta\lambda_{1},&\lambda<\lambda_{p}\\ \Delta\lambda_{2},&\lambda\geq\lambda_{p}\end{cases}$ (3)
$PSD_{i}(\lambda)=\frac{g(\lambda,\lambda_{p},\Delta\lambda)+k_{1}g^{k_{2}}(\lambda,\lambda_{p},\Delta\lambda)}{1+k_{1}}$ (4)
Where $\lambda_{p}$ is the peak wavelength, $\Delta\lambda_{1}$ and $\Delta\lambda_{2}$ are the left/right half spectral widths, and $k_{1}$, $k_{2}$ are characteristic shape parameters for color channel $i$.
@@
$p_{i,j}^{(k)}=\frac{1}{3}\int_{\lambda_{l}^{i}}^{\lambda_{h}^{i}}PSD(\lambda)F_{j}(\lambda)S_{i}h^{(k)}d\lambda;i,j\in\{1,2,3\}$ (5)
In the context provided, $i$ and $j$ correspond to the colors of the LED at the transmitter and the photodetector at the receiver, respectively. The values of $i$ and $j$ range from $1$ to $3$, representing the colors red, green, and blue. The variables $\lambda_{l}^{i}$ and $\lambda_{h}^{i}$ denote the lower and upper wavelengths of the spectrum for the LED color $i$, respectively. The variable $F_{j}$ represents the gain associated with the $j$-colored optical filter. The variable $S_{i}$ denotes the CSK symbol value for color $i$, and finnaly $h$ is the VLC channel attenuation. The matrix $\boldsymbol{p}$ is a $3\times 3$ channel matrix, where the $(i,j)$-th element is determined by Eq. 5.
$\boldsymbol{p}=\begin{bmatrix}p_{1,1}&p_{2,1}&p_{3,1}\\ p_{1,2}&p_{2,2}&p_{3,2}\\ p_{1,3}&p_{2,3}&p_{3,3}\end{bmatrix}$ (6)
@@
$\operatorname{SINR}=\frac{\sum_{i,j}p_{i,j}\delta_{ij}}{\sum_{i,j}p_{i,j}-\sum_{i,j}p_{i,j}\delta_{ij}+N_{0}}$ (7)
Here, $\delta_{ij}$ is the Kronecker delta function and $N_{0}$ is the power of AWGN noise respectively. Capacity calculation over SINR can be done according to Eq.8. After the CSK demodulation is performed on the received signal, the throughput can be calculated using Eq.9.
$C=B\log_{2}(1+\text{SINR})$ (8)
$T=C(1-\text{BER})$ (9)
The open-source program shared in [32] can be used to calculate the illumination metrics such as CRI [8], CCT [7], and luminous flux. The irradiance data corresponding to the LED’s wavelength is input into the program, which then uses it to calculate the illumination parameters. The CRI, CCT, and illumination values should be calculated over the long-term average of the irradiance value because the brightness of the light emitted from the R, G, and B colors varies with communication.
4 Numerical Results
In this section, we simulated the environment to evaluate the performance of the proposed NOMA-CSK system for VLC networks. The variables associated with the VLC channel, along with simulation parameters, are detailed in Table 2. Our simulation setup includes a NOMA-CSK communication system with two receivers ($U_{1}$ and $U_{2}$), each equipped with three photodetectors (PDs) positioned perpendicular to the ceiling.